OpenStudy (anonymous):

HELPPP!! If sin thea 4/5 = and thea is in quadrant 2, the value of cot thea is _____. If necessary, use the slash mark ( / ) for a fraction bar.

OpenStudy (raden):

use the identity : cos^2 a = 1 - sin^2 a

OpenStudy (raden):

so, cos^2 a = 1 - (4/5)^2 cos^2 a = 1 - 16/25 cos^2 a = 25/25 - 16/25 cos^2 a = 9/25 cos a = sqrt(9/25) = 3/5

OpenStudy (raden):

because the a in the 2nd quadrant, it must be negative for cos therefore, cos a = - 3/5

OpenStudy (anonymous):

it was -3/4 thnks tho